### v2

@@ Appendix A Proofs
[…]
@@
[…]
@@
where
$\tilde{h}(x,y)=\begin{cases}\hat{H}_{j}(x,y)-H(x,y)\text{ if }y\leq x,\\ H(x,y)-\hat{H}_{j}(x,y)\text{ if }x<y.\\ \end{cases}$
@@
[…]
@@
Now, let $\varepsilon=C_{1}n^{-c_{1}}/\sqrt{\log p}$ and take $\delta=C_{\delta}n^{-r}$ for some $r>1/7$. Recall that, by assumption, $\log p\leq C_{3}n^{a}$ for some $a<1/7$. With this choice of $r$, we can choose $C_{\delta}$ sufficiently large and $c_{1}$ sufficiently small such that $38\sqrt{\delta\log(1/\delta)}<\varepsilon/$ 3. Using (10), (12), we have that
@@
Suppose that Assumption 1 holds and there exist constants $C_{1}>0$ and $40<\gamma<1/$ 4 such that $(1/q)\log^{2}p\leq C_{1}n^{-\gamma}$ and $\max\{q\log^{1/2}p,\log^{3/2}p,\sqrt{q}\log^{7/2}(pn)\}\leq C_{1}n^{1/2-\gamma}$. Then there exist constants $c,C>0$ depending only on $\gamma$ and $C_{1}$ such that, under the null hypothesis $H_{0}$,
$\sup_{t\in\mathbb{R}}\left\lvert{\mathrm{P}}\left(T_{0}\leq t\right)-{\mathrm{P}}\left(Z_{0}\leq t\right)\right\rvert\leq Cn^{-c}\leavevmode\nobreak\ .$ (13)
@@
Lemma 3 is similar to Theorem E.1 in Chernozhukov, Chetverikov, and Kato (2019), but it differs from the latter in two ways: the mixing rate does not appear in our bound and we do not need to assume that the small blocks grow with the sample size. This difference stems from our derivations directly imposing the 1-dependence of $\{W_{j,i}\}_{i=1}^{n}$ rather than a general mixing condition.
@@
$-Cn^{-c}\leq{\mathrm{P}}(T_{0}\leq t)-{\mathrm{P}}\left(\max_{1\leq j\leq p}\sqrt{\frac{mq}{n}}\,V_{j}\leq t\right)\leq Cn^{-c}$ (14)
[…]
@@
In steps 1–3, we show that each of the terms in (i)–(iii) is bounded in the supremum norm (w.r.t. $t$) by $Cn^{-c}$. To show the first and second inequality in (14), we use $\star$=-⋆ = - and =$\star$=+⋆ = +, respectively.
@@
Since $\{A_{j,k}\}_{k=1}^{m}$ are independent, we may apply Corollary 2.1 in Chernozhukov, Chetverikov, and Kato (2013) (note that the covariance matrix of $\sqrt{mq/n}\,V$ is the same as that of $n^{-1/2}\sum_{k=1}^{m}A_{k,j}$). To this end, we seek to verify the conditions of the corollary applied to our case. Observe that
@@
$\bar{\sigma}^{2}(q)=\max_{1\leq j\leq p}\max_{I}\text{Var}\left(\frac{1}{\sqrt{q}}\sum_{i\in I}W_{ij}\right)\quad\text{ and }\quad\underline{\sigma}^{2}(q)=\min_{1\leq j\leq p}\min_{I}\text{Var}\left(\frac{1}{\sqrt{q}}\sum_{i\in I}W_{ij}\right)\leavevmode\nobreak\ ,$
where the $\max_{I}$ and $\min_{I}$ are taken over all $I\subset\{1,\dots,n\}$ of the form $I=\{i+1,\dots,i+q\}$. Angus (1995) shows that $\operatorname*{\mathbb{E}}(W_{j,i})=0$, ${\rm Var}(W_{j,i})=1/2$, and $Cov(W_{j,i},W_{j,i+k})=-1/20$ for $k=1$ (00 otherwise). Therefore,
$\text{Var}\left(\frac{1}{\sqrt{q}}\sum_{i\in I}W_{i,j}\right)$ $=\frac{1}{q}\left(\sum_{i\in I}\text{Var}(W_{i,j})+2\sum_{i=1}^{q-1}\,\text{Cov}(W_{j,i},W_{j,i+1})\right)=\frac{2}{5}+\frac{1}{10q}\leavevmode\nobreak\ ,$
@@
$\frac{2}{5}\frac{1}{4}\leq\frac{\underline{\sigma}^{2}(q)}{4}\leq{\rm Var}\left(\frac{A_{j,k}}{\sqrt{n/m}}\right)\leq\bar{\sigma}^{2}(q)\leq\frac{1}{2},\quad 1\leq j\leq p\leavevmode\nobreak\ ,$
and $\left\lvert A_{j,k}/\sqrt{n/m}\right\rvert\leq 2\sqrt{q}\,$ a.s. so the conditions of Corollary 2.1 (i) in Chernozhukov, Chetverikov, and Kato (2013) are verified with $B_{n}=2\sqrt{q}$, which leads to the assertion of this step (note that $C^{-1}n^{c}\leq n/(4q)\leq m$).
@@
$\frac{2}{5}\leq\underline{\sigma}^{2}(q)\leq{\rm Var}(V_{j})\leq\bar{\sigma}^{2}(q)\leq\frac{1}{2},\quad 1\leq j\leq p\leavevmode\nobreak\ ,$
the desired result follows by Corollary 1 in Chernozhukov, Chetverikov, and Kato (2015) restated in Lemma 5. We apply this result with $\varepsilon=C/(2n^{c}\sqrt{\log p})(mq/n)^{-1/2}$ and conclude that
@@
Let $\varepsilon_{1},\dots,\varepsilon_{m}$ be independent standard normal random variables, independent of the data $\mathbb{D}$. Suppose that there exist constants $C_{1}>0$ and $20<\gamma<1/$ 2 such that $q\log^{5/2}p\leq C_{1}\,n^{1/2-\gamma}$. Then, under Assumption 1 and the null hypothesis $H_{0}$, there exist constants $c,c^{\prime},C,C^{\prime}>0$ depending only on $\gamma$ and $C_{1}$ such that, with probability larger than $1-Cn^{-c}$,
@@
Here $c,c^{\prime},C,C^{\prime}$ denote generic positive constants depending only on $\gamma$; their values may change from place to place. By Theorem 2 in Chernozhukov, Chetverikov, and Kato (2015), the left side in the conclusion of the lemma is bounded by $CD^{1/3}\max\{1,\log(p/D)\}^{2/3}$, where
$D=\max_{1\leq j,j^{\prime}\leq p}\,\left\lvert\frac{1}{mq}\sum_{k=1}^{m}\bigl{(}A_{j,k}A_{j^{\prime},k}-\operatorname*{\mathbb{E}}[A_{j,k}A_{j^{\prime},k}]\bigr{)}\right\rvert\leavevmode\nobreak\ .$
@@
P(${\mathrm{P}}\left({\mathrm{P}}\left(\left\lvert\hat{T}^{B}-T^{B}_{0}\right\rvert>\zeta_{n}|\mathbb{D}\right)>Cn^{-c}\right)\leq Cn^{-c}\leavevmode\nobreak\ .$ (17)
@@
$\left\lvert\hat{T}-T_{0}\right\rvert$ $\leq\max_{1\leq j\leq p}\left\lvert\hat{\xi}_{j}-\frac{1}{\sqrt{n}}\sum_{i=1}^{n-1}W_{j,i}\right\rvert\leq\max_{1\leq j\leq p}\left\lvert r_{j,n}\right\rvert=\Delta_{1}\leavevmode\nobreak\ .$
@@
We first show (21). By Theorem 3 in Chernozhukov, Chetverikov, and Kato (2015), for any $t\in\mathbb{R}$ and any $\epsilon>0$,
@@
with $22/5\leq E[A_{j,k}^{2}]/q\leq 1/$ 2, as shown in the proof of Lemma 3. Therefore,
@@
${\mathrm{P}}(Z_{0}\leq c(\gamma+20\zeta_{n}\sqrt{\log p})+\zeta_{n})$ $\leq{\mathrm{P}}(Z_{0}\leq c(\gamma+20\zeta_{n}\sqrt{\log p}))+2\zeta_{n}(\sqrt{2\log p}+1)/\sigma_{V}$
$\leq{\mathrm{P}}(Z_{0}\leq c(\gamma+20\zeta_{n}\sqrt{\log p}))+8\zeta_{n}\sqrt{\log p}\cdot\frac{5}{2}$
$\leq 1-\gamma-20\zeta_{n}\sqrt{\log p}+20\zeta_{n}\sqrt{\log p})$
@@
×$\quad\quad\times{\mathrm{P}}\left(\hat{c}(\alpha)<c(\alpha+Cn^{-c}+20\zeta_{n}\sqrt{\log p})\right)$
@@
$\leq{\mathrm{P}}\left(\hat{T}\leq r_{n}\right)+{\mathrm{P}}\left(\hat{c}(\alpha)\geq v_{n}\right)+{\mathrm{P}}\left(v_{n}\geq r_{n}\right)\leavevmode\nobreak\ .$
@@
${\mathrm{P}}\left(\hat{T}\leq r_{n}\right)$ $=o(1)\leavevmode\nobreak\ .$ (23)
@@
$\operatorname*{\mathbb{E}}\left[\hat{T}^{B}|\mathbb{D}\right]\leq 2\sqrt{2q\log p}\leavevmode\nobreak\ .$
@@
${\mathrm{P}}\left\{\left.\max_{1\leq j\leq p}\hat{T}_{j}^{B}>\operatorname*{\mathbb{E}}\left[\left.\max_{1\leq j\leq p}\hat{T}_{j}^{B}\;\right|\;\mathbb{D}\right]+\lambda\;\right|\;\mathbb{D}\right\}\leq\alpha\leavevmode\nobreak\ .$
By the definition of $\hat{c}(\alpha)$, together with the anti-concentration inequality (Chernozhukov, Chetverikov, and Kato, 2015, Corollary 1), we have that
${\mathrm{P}}\left(\left.\max_{1\leq j\leq p}\hat{T}_{j}^{B}>\hat{c}(\alpha)\;\right|\;\mathbb{D}\right)=\alpha\leavevmode\nobreak\ .$
@@
We now establish (23). Note that under the alternative, at least one of the individual hypotheses $H_{0,j}:\,Y_{j}\bot X$, $1\leq j\leq p$, fails. Let $j^{*}$ be the index of one of these hypotheses, so that $\xi_{j^{*}}>0$ (since $\xi_{j}$ is nonnegative for any $j$). Define $Z_{n}\coloneqq\sqrt{n}(\hat{\xi}_{j^{*}}-\xi_{j^{*}})$. It follows from Kroll (2024) that $Z_{n}=O_{P}(1)$. Further, observe that
$\hat{T}=\sqrt{n}\max_{1\leq j\leq p}\,\hat{\xi}_{j}\geq\sqrt{n}\hat{\xi}_{j^{*}}=Z_{n}+\sqrt{n}\xi_{j^{*}}\leavevmode\nobreak\ .$
For any sequence $r_{n}=o(n^{1/2})$, we have that
${\mathrm{P}}\left(\hat{T}>r_{n}\right)$ $\geq{\mathrm{P}}\left(Z_{n}+\sqrt{n}\xi_{j^{*}}>r_{n}\right)$
$={\mathrm{P}}\left(Z_{n}+\sqrt{n}\xi_{j^{*}}>r_{n}\;\big{|}\;|Z_{n}|\leq\sqrt{n}\xi_{j^{*}}/2\right)\,{\mathrm{P}}\left(\left\lvert Z_{n}\right\rvert\leq\sqrt{n}\xi_{j^{*}}/2\right)$
$\quad+{\mathrm{P}}\left(Z_{n}+\sqrt{n}\xi_{j^{*}}>r_{n}\;\big{|}\;|Z_{n}|>\sqrt{n}\xi_{j^{*}}/2\right)\,{\mathrm{P}}\left(\left\lvert Z_{n}\right\rvert>\sqrt{n}\xi_{j^{*}}/2\right)$
$\geq{\mathrm{P}}\left(\sqrt{n}\xi_{j^{*}}/2>r_{n}\right)\left(1-o(1)\right)+o(1)$
$=1-o(1).$

### v3

@@ Appendix A Proofs
Part (i). Note that $\hat{F}_{Y_{j}}(Y_{j,(i)})=\hat{F}_{U_{j}}(U_{j,i})$, where $U_{j,i}\coloneqq F_{Y_{j}}(Y_{j,(i)})$ and $\hat{F}_{U_{j}}(\cdot)$ is the empirical CDF of $\{U_{j,i}\}_{i=1}^{n}$. Recall that under the null hypothesis of independence and Assumption 1, for each $j$, $\{U_{j,i}\}_{i=1}^{n}$ are independent random variables distributed uniformly on $[0,1]$. Following Angus (1995), we note that
@@
$\int_{0}^{1}\int_{0}^{1}\left(\left\lvert\hat{F}_{U_{j}}(y)-\hat{F}_{U_{j}}(x)\right\rvert-|y-x|\right)dydx=\frac{1}{n}\sum_{i=1}^{n}2U_{j,i}(U_{j,i}-1)-\frac{1}{3}.$
@@
$|I_{j,n}|\leq 8\left\lVert\hat{B}_{j}(x)-B_{j}(x)\right\rVert_{\infty}+\left\lvert\int_{0}^{1}\int_{0}^{1}B_{j}(x)d\tilde{h}(x,y)\right\rvert+\left\lvert\int_{0}^{1}\int_{0}^{1}B_{j}(y)d\tilde{h}(x,y)\right\rvert$, (11)
where
@@
$\left|\int_{0}^{1}\int_{0}^{1}B_{j}(x)d\tilde{h}(x,y)\right|$ $=\Big|\int_{0}^{1}\int_{0}^{1}\big(B_{j}(x)-B^{\delta}_{j}(x)\big)d\tilde{h}(x,y)+\int_{0}^{1}\int_{0}^{1}B^{\delta}_{j}(x)d\tilde{h}(x,y)\Big|$
@@
Now, let $\varepsilon=C_{1}n^{-c_{1}}/\sqrt{\log p}$ and take $\delta=C_{\delta}n^{-r}$ for some $r>1/7$. Recall that, by assumption, $\log p\leq C_{3}n^{a}$ for some $a<1/7$. With this choice of $r$, we can choose $C_{\delta}$ sufficiently large and $c_{1}$ sufficiently small such that $8\sqrt{\delta\log(1/\delta)}<\varepsilon/3$. Using (10), (12), we have that
@@
Suppose that Assumption 1 holds and there exist constants $C_{1}>0$ and $0<\gamma<1/4$ such that $(1/q)\log^{2}p\leq C_{1}n^{-\gamma}$ and $\max\{q\log^{1/2}p,\log^{3/2}p,\sqrt{q}\log^{7/2}(pn)\}\leq C_{1}n^{1/2-\gamma}$. Then there exist constants $c,C>0$ depending only on $\gamma$ and $C_{1}$ such that, under the null hypothesis $H_{0}$,
$\sup_{t\in\mathbb{R}}\left\lvert{\mathrm{P}}\left(T_{0}\leq t\right)-{\mathrm{P}}\left(Z_{0}\leq t\right)\right\rvert\leq Cn^{-c}\penalty 10000\ .$ (13)
@@
Lemma 3 is similar to Theorem E.1 in Chernozhukov et al. (2019), but it differs from the latter in two ways: the mixing rate does not appear in our bound and we do not need to assume that the small blocks grow with the sample size. This difference stems from our derivations directly imposing the 1-dependence of $\{W_{j,i}\}_{i=1}^{n}$ rather than a general mixing condition.
@@
$Cn^{-c}\leq{\mathrm{P}}(T_{0}\leq t)-{\mathrm{P}}\left(\max_{1\leq j\leq p}\sqrt{\frac{mq}{n}}\,V_{j}\leq t\right)\leq Cn^{-c}$ (14)
for all $t\in\mathbb{R}$. For $\star\in\{+,-\}$, consider the following decomposition
@@
In steps 1–3, we show that each of the terms in (i)–(iii) is bounded in the supremum norm (w.r.t. $t$) by $Cn^{-c}$. To show the first and second inequality in (14), we use $\star=$- and $\star=+$, respectively.
@@
Since $\{A_{j,k}\}_{k=1}^{m}$ are independent, we may apply Corollary 2.1 in Chernozhukov et al. (2013) (note that the covariance matrix of $\sqrt{mq/n}\,V$ is the same as that of $n^{-1/2}\sum_{k=1}^{m}A_{k,j}$). To this end, we seek to verify the conditions of the corollary applied to our case. Observe that
@@
$\bar{\sigma}^{2}(q)=\max_{1\leq j\leq p}\max_{I}\text{Var}\left(\frac{1}{\sqrt{q}}\sum_{i\in I}W_{ij}\right)\quad\text{ and }\quad\underline{\sigma}^{2}(q)=\min_{1\leq j\leq p}\min_{I}\text{Var}\left(\frac{1}{\sqrt{q}}\sum_{i\in I}W_{ij}\right)\penalty 10000\ ,$
where the $\max_{I}$ and $\min_{I}$ are taken over all $I\subset\{1,\dots,n\}$ of the form $I=\{i+1,\dots,i+q\}$. Angus (1995) shows that $\operatorname*{\mathbb{E}}(W_{j,i})=0$, ${\rm Var}(W_{j,i})=1/2$, and $Cov(W_{j,i},W_{j,i+k})=-1/20$ for $k=1$ (0 otherwise). Therefore,
$\text{Var}\left(\frac{1}{\sqrt{q}}\sum_{i\in I}W_{i,j}\right)$ $=\frac{1}{q}\left(\sum_{i\in I}\text{Var}(W_{i,j})+2\sum_{i=1}^{q-1}\,\text{Cov}(W_{j,i},W_{j,i+1})\right)=\frac{2}{5}+\frac{1}{10q}\penalty 10000\ ,$
@@
$\frac{2}{5}\frac{1}{4}\leq\frac{\underline{\sigma}^{2}(q)}{4}\leq{\rm Var}\left(\frac{A_{j,k}}{\sqrt{n/m}}\right)\leq\bar{\sigma}^{2}(q)\leq\frac{1}{2},\quad 1\leq j\leq p\penalty 10000\ ,$
and $\left\lvert A_{j,k}/\sqrt{n/m}\right\rvert\leq 2\sqrt{q}\,$ a.s. so the conditions of Corollary 2.1 (i) in Chernozhukov et al. (2013) are verified with $B_{n}=2\sqrt{q}$, which leads to the assertion of this step (note that $C^{-1}n^{c}\leq n/(4q)\leq m$).
@@
$\frac{2}{5}\leq\underline{\sigma}^{2}(q)\leq{\rm Var}(V_{j})\leq\bar{\sigma}^{2}(q)\leq\frac{1}{2},\quad 1\leq j\leq p\penalty 10000\ ,$
the desired result follows by Corollary 1 in Chernozhukov et al. (2015) restated in Lemma 5. We apply this result with $\varepsilon=C/(2n^{c}\sqrt{\log p})(mq/n)^{-1/2}$ and conclude that
@@
Let $\varepsilon_{1},\dots,\varepsilon_{m}$ be independent standard normal random variables, independent of the data $\mathbb{D}$. Suppose that there exist constants $C_{1}>0$ and $0<\gamma<1/2$ such that $q\log^{5/2}p\leq C_{1}\,n^{1/2-\gamma}$. Then, under Assumption 1 and the null hypothesis $H_{0}$, there exist constants $c,c^{\prime},C,C^{\prime}>0$ depending only on $\gamma$ and $C_{1}$ such that, with probability larger than $1-Cn^{-c}$,
@@
Here $c,c^{\prime},C,C^{\prime}$ denote generic positive constants depending only on $\gamma$; their values may change from place to place. By Theorem 2 in Chernozhukov et al. (2015), the left side in the conclusion of the lemma is bounded by $CD^{1/3}\max\{1,\log(p/D)\}^{2/3}$, where
$D=\max_{1\leq j,j^{\prime}\leq p}\,\left\lvert\frac{1}{mq}\sum_{k=1}^{m}\bigl(A_{j,k}A_{j^{\prime},k}-\operatorname*{\mathbb{E}}[A_{j,k}A_{j^{\prime},k}]\bigr)\right\rvert\penalty 10000\$.
@@
${\mathrm{P}}\left({\mathrm{P}}\left(\left\lvert\hat{T}^{B}-T^{B}_{0}\right\rvert>\zeta_{n}|\mathbb{D}\right)>Cn^{-c}\right)\leq Cn^{-c}\penalty 10000\ .$ (17)
@@
$\left\lvert\hat{T}-T_{0}\right\rvert$ $\leq\max_{1\leq j\leq p}\left\lvert\hat{\xi}_{j}-\frac{1}{\sqrt{n}}\sum_{i=1}^{n-1}W_{j,i}\right\rvert\leq\max_{1\leq j\leq p}\left\lvert r_{j,n}\right\rvert=\Delta_{1}\penalty 10000\ .$
@@
We first show (21). By Theorem 3 in Chernozhukov et al. (2015), for any $t\in\mathbb{R}$ and any $\epsilon>0$,
@@
with $2/5\leq E[A_{j,k}^{2}]/q\leq 1/2$, as shown in the proof of Lemma 3. Therefore,
@@
${\mathrm{P}}(Z_{0}\leq c(\gamma+20\zeta_{n}\sqrt{\log p})+\zeta_{n})$ $\leq{\mathrm{P}}\left(Z_{0}\leq c(\gamma+20\zeta_{n}\sqrt{\log p})\right)+2\zeta_{n}(\sqrt{2\log p}+1)/\sigma_{V}$
$\leq{\mathrm{P}}\left(Z_{0}\leq c(\gamma+20\zeta_{n}\sqrt{\log p})\right)+8\zeta_{n}\sqrt{\log p}\cdot\frac{5}{2}$
$\leq 1-\gamma-20\zeta_{n}\sqrt{\log p}+20\zeta_{n}\sqrt{\log p}$
@@
$\quad\quad\times{\mathrm{P}}\left(\hat{c}(\alpha)<c(\alpha+Cn^{-c}+20\zeta_{n}\sqrt{\log p})\right)$
@@
$\leq{\mathrm{P}}\left(\hat{T}\leq r_{n}\right)+{\mathrm{P}}\left(\hat{c}(\alpha)\geq v_{n}\right)+{\mathrm{P}}\left(v_{n}\geq r_{n}\right)\penalty 10000\ .$
@@
${\mathrm{P}}\left(\hat{T}\leq r_{n}\right)$ $=o(1)\penalty 10000\ .$ (23)
@@
$\operatorname*{\mathbb{E}}\left[\hat{T}^{B}|\mathbb{D}\right]\leq 2\sqrt{2q\log p}\penalty 10000\ .$
@@
[…]
By the definition of $\hat{c}(\alpha)$, together with the anti-concentration inequality (Chernozhukov et al., 2015, Corollary 1), we have that
${\mathrm{P}}\left(\left.\max_{1\leq j\leq p}\hat{T}_{j}^{B}>\hat{c}(\alpha)\;\right|\;\mathbb{D}\right)=\alpha\penalty 10000\ .$
@@
We now establish (23). Note that under the alternative, at least one of the individual hypotheses $H_{0,j}:\,Y_{j}\bot X$, $1\leq j\leq p$, fails. Let $j^{*}$ be the index of one of these hypotheses, so that $\xi_{j^{*}}>0$ (since $\xi_{j}$ is nonnegative for any $j$). Observe that $\hat{T}=\sqrt{n}\max_{1\leq j\leq p}\,\hat{\xi}_{j}\geq\sqrt{n}\,\hat{\xi}_{j^{*}}$. Since $r_{n}=o(n^{1/2})$ and $\xi_{j^{*}}>0$, we have that $r_{n}/\sqrt{n}-\xi_{j^{*}}<-\xi_{j^{*}}/2$ for sufficiently large $n$. Then, strong consistency of $\hat{\xi}_{j^{*}}$ (Chatterjee, 2021, Theorem 1.1) implies
${\mathrm{P}}\left(\hat{T}>r_{n}\right)$ $\geq{\mathrm{P}}\left(\hat{\xi}_{j^{*}}-\xi_{j^{*}}>\frac{r_{n}}{\sqrt{n}}-\xi_{j^{*}}\right)\geq{\mathrm{P}}\left(\hat{\xi}_{j^{*}}-\xi_{j^{*}}>-\frac{\xi_{j^{*}}}{2}\right)=1-o(1)\penalty 10000\ .$